ZSOLT TUZA

List of Publications

1. Induced subtrees in graphs of large chromatic number. Discrete Mathematics 30 (3), 1980, 235–244. (with A. Gyárfás and E. Szemerédi)

2. Ensembles d’articulation d’un graphe γ-critique. Discrete Mathematics 30 (3), 1980, 245–248. (with J. Lehel)

3. τ-critical hypergraphs. “ Ninth Winter School on Abstract Analysis ” (Z. Frolík, ed.), Math. Inst. Czechosl. Acad. Sci., Praha, 1981, 181–184.

4. Worst-case behavior of string-searching algorithms. Journal of Statistical Planning and Inference 6 (1), 1982, 99–103.

5. Upper bound on the order of τ-critical hypergraphs. Journal of Combinatorial Theory, Ser. B 33 (2), 1982, 161–165. (with A. Gyárfás and J. Lehel)

6. Triangle-free partial graphs and edge covering theorems. Discrete Mathematics 39 (1), 1982, 59–65. (with J. Lehel)

7. Upper bound for transversals of tripartite hypergraphs. Periodica Mathematica Hungarica 13 (4), 1982, 321–323. (with E. Szemerédi)

8. Ryser’s conjecture on transversals of r-partite hypergraphs. Ars Combinatoria 16B, 1983, 201–209.

9. The method of set-pairs for extremal problems in hypergraphs. “ Finite and Infinite Sets ” (A. Hajnal et al., eds.), North-Holland, 1984, 749–762.

10. Covering of graphs by complete bipartite subgraphs; complexity of 0–1 matrices. Combinatorica 4 (1), 1984, 111–116.

11. Helly-type hypergraphs and Sperner families. European Journal of Combinatorics 5 (2), 1984, 185–187.

12. The structure of rectangle families dividing the plane into maximum number of atoms. Discrete Mathematics 52 (2–3), 1984, 177–198. (with A. Gyárfás and J. Lehel)

13. Computing N with minimal number of additions. “ Theory of Algorithms ” (L. Lovász and E. Szemerédi, eds.), North-Holland, 1985, 371–373. (with I. Z. Ruzsa)

14. On color critical graphs. Journal of Combinatorial Theory, Ser. B 38 (3), 1985, 204–213. (with V. Rödl)

15. Hypergraphs without a large star. Discrete Mathematics 55 (3), 1985, 317–321. (with Z. Füredi)

16. Critical hypergraphs and intersecting set-pair systems. Journal of Combinatorial Theory, Ser. B 39 (2), 1985, 134–145.

17. How many atoms can be defined by boxes? Combinatorica 5 (3), 1985, 193–204. (with A. Gyárfás and J. Lehel)

18. Saturated graphs with minimal number of edges. Journal of Graph Theory 10 (2), 1986, 203–210. (with L. Kászonyi)

19. Neighborhood perfect graphs. Discrete Mathematics 61 (1), 1986, 93–101. (with J. Lehel)

20. A generalization of saturated graphs for finite languages. “ Proceedings of IMYCS ’86 ” (J. Demetrovics and J. Kelemen, eds.), MTA SZTAKI Studies 185, Comp. Autom. Inst. Hungar. Acad. Sci., Budapest, 1986, 287–293.

21. On the order of vertex sets meeting all edges of a 3-partite hypergraph. Ars Combinatoria 24A, 1987, 59–63.

22. Linear upper bounds for local Ramsey numbers. Graphs and Combinatorics 3 (1), 1987, 67–73. (with M. Truszczyński)

23. Inequalities for two set systems with prescribed intersections. Graphs and Combinatorics 3 (1), 1987, 75–80.

24. On two intersecting set systems and k-continuous Boolean functions. Discrete Applied Mathematics 16 (2), 1987, 183–185.

25. Maximum bipartite subgraphs of Kneser graphs. Graphs and Combinatorics 3 (2), 1987, 191–199. (with S. Poljak)

26. Decompositions of graphs into complete subgraphs of given order. Studia Scientiarum Mathematicarum Hungarica 22 (3–4), 1987, 315–320. (with E. Győri)

27. Ramsey numbers for local colorings. Graphs and Combinatorics 3 (3), 1987, 267–277. (with A. Gyárfás, J. Lehel, and R. H. Schelp)

28. Another extremal property of some Turán graphs. European Journal of Combinatorics 8 (1), 1987, 27–28. (with C. S. Edwards and P. D. Johnson)

29. An upper bound on the Ramsey number of trees. Discrete Mathematics 66 (3), 1987, 309–310. (with A. Gyárfás)

30. On the context-free production complexity of finite languages. Discrete Applied Mathematics 18 (3), 1987, 293–304.

31. Local k-colorings of graphs and hypergraphs. Journal of Combinatorial Theory, Ser. B 43 (2), 1987, 127–139. (with A. Gyárfás, J. Lehel, J. Nešetřil, V. Rödl, and R. H. Schelp)

32. Some combinatorial problems concerning finite languages. “ Trends, Techniques and Problems in Theoretical Computer Science ” (A. Kelemenová, J. Kelemen, eds.), Lecture Notes in Computer Science 281, Springer-Verlag, 1987, 51–58.

33. Menger’s theorem and short paths. Journal of Combinatorial Mathematics and Combinatorial Computing 2, 1987, 235–253. (with R. J. Faudree, M. S. Jacobson, E. T. Ordman, and R. H. Schelp)

34. P_{3}-factors and covering cycles in graphs of minimum degree n/3. “ Combinatorics ” (A. Hajnal et al., eds.), North-Holland, 1988, 213–220. (with H. Enomoto and A. Kaneko)

35. Extremal problems on saturated graphs and hypergraphs. Ars Combinatoria 25B, 1988, 105–113.

36. Isomorphism of two infinite-chromatic triangle-free graphs. Discrete Mathematics 69 (1), 1988, 95.

37. Clumsy packing of dominoes. Discrete Mathematics 71 (1), 1988, 33–46. (with A. Gyárfás and J. Lehel)

38. Graphs in which all spanning subgraphs with r edges fewer are isomorphic. Congressus Numerantium 67, 1988, 45–58. (with P. Duchet and P. D. Vestergaard)

39. Planar division patterns and map generating D0L systems. “ Second Conference on Automata, Languages and Programming Systems ” (F. Gécseg and I. Peák, eds.), Karl Marx Univ. of Economics Press, DM 88–4, Budapest, 1988, 229–233.

40. Decompositions of partially ordered sets into chains and antichains of given size. Order 5 (3), 1988, 245–255. (with Y. Caro)

41. Vertex partitions of graphs into k-stable and k-complete subgraphs. Ars Combinatoria 27, 1989, 61–62.

42. Langford strings are square-free. International Journal of Computer Mathematics 29 (2– 4), 1989, 75–78.

43. Intersection properties and extremal problems for set systems. “ Irregularities of Partitions ” (G. Halász and V. T. Sós, eds.), Algorithms and Combinatorics 8, Springer-Verlag, 1989, 141–151.

44. On the maximum number of qualitatively independent partitions. Journal of Combinatorial Theory, Ser. A 51 (1), 1989, 111–116. (with S. Poljak)

45. Minimum number of elements representing a set system of given rank. Journal of Combinatorial Theory, Ser. A 52 (1), 1989, 84–89.

46. Radius, diameter and minimum degree. Journal of Combinatorial Theory, Ser. B 47 (1), 1989, 73–79. (with P. Erdős, J. Pach, and R. Pollack)

47. Induced matchings in bipartite graphs. Discrete Mathematics 78 (1–2), 1989, 83–87. (with R. J. Faudree, A. Gyárfás, and R. H. Schelp)

48. An application of set-pair systems for multitransversals. Acta Universitatis Carolinae, Mathematica et Physica 30 (2), 1989, 37–39. (with Y. Caro)

49. Sub-Ramsey numbers for arithmetic progressions. Graphs and Combinatorics 5 (4), 1989, 303–306. (with N. Alon and Y. Caro)

50. C_{4}-saturated graphs of minimum size. Acta Universitatis Carolinae, Mathematica et Physica 30 (2), 1989, 161–167.

51. An exact value for the path-chromatic number of a graph. Le Matematiche 44 (2), 1989, 345–350. (with F. Harary)

52. Encoding the vertices of a graph with binary edge labels. “ Sequences—Combinatorics, Compression, Security, and Transmission ” (R. M. Capocelli, ed.), Springer-Verlag, New York, 1990, 287–299.

53. Covering all cliques of a graph. Discrete Mathematics 86 (1–3), 1990, 117–126.

reprinted in: “ Topics on Domination,” Annals of Discrete Mathematics 48, 1991.

54. Rainbow Hamiltonian paths and canonically colored subgraphs in infinite complete graphs. Mathematica Pannonica 1 (1), 1990, 5–13. (with P. Erdős)

55. Contractions and minimal k-colorability. Graphs and Combinatorics 6 (1), 1990, 51–59.

56. A characterization of graphs without long induced paths. Journal of Graph Theory 14 (4), 1990, 455–464. (with G. Bacsó)

57. The strong chromatic index of graphs. Ars Combinatoria 29B, 1990, 205–211. (with R. J. Faudree, A. Gyárfás, and R. H. Schelp)

58. The maximum number of edges in 2K_{2}-free graphs of bounded degree. Discrete Mathematics 81 (2), 1990, 129–135. (with F. R. K. Chung, A. Gyárfás, and W. T. Trotter)

59. Polynomial algorithms for finding cycles and paths in bipartite tournaments. SIAM Journal on Discrete Mathematics 3 (4), 1990, 537–543. (with Y. Manoussakis)

60. Self-generating hexagonal cell division patterns. Geometriae Dedicata 34 (1), 1990, 19–34. (with A. Lindenmayer and J. Malitz)

61. Periodic string division generated by deterministic L systems. Information Processing Letters 34 (5), 1990, 229–234.

62. Extremal problems on finite set systems. “ International Conference on Discrete Mathematics ” (H. Sachs, ed.), Technische Hochschule Ilmenau Press, 1990, 70–73.

63. Matchings and coverings in regular uniform hypergraphs. Ars Combinatoria 29C, 1990, 122–129.

64. A conjecture on triangles of graphs. Graphs and Combinatorics 6 (4), 1990, 373–380.

65. Exponentially many distinguishable cycles in graphs. Journal of Combinatorics, Information and System Sciences 15 (1–4), 1990, 281–285.

66. A coloring problem related to the Erdős–Faber–Lovász conjecture. Journal of Combinatorial Theory, Ser. B 50 (2), 1990, 321–322; 51 (2), 1991, 329. (with P. Horák)

67. Dominating cliques in P_{5}-free graphs. Periodica Mathematica Hungarica 21 (4), 1990, 303–308. (with G. Bacsó)

68. Problems and results on graph and hypergraph colorings. Le Matematiche 45 (1), 1990, 219-238.

69. Asymptotic results on saturated graphs. Discrete Mathematics 87 (3), 1991, 309–314. (with M. Truszczyński)

70. Languages defined by distance constraints. Computers and Artificial Intelligence 10 (2), 1991, 167–172.

71. Improved lower bounds on k-independence. Journal of Graph Theory 15 (1), 1991, 99–107. (with Y. Caro)

72. Clique-transversal sets of line graphs and complements of line graphs. Discrete Mathematics 88 (1), 1991, 11–20. (with Th. Andreae and M. Schughart)

73. Hypergraph coverings and local colorings. Journal of Combinatorial Theory, Ser. B 52 (1), 1991, 79–85. (with Y. Caro)

74. A periodic division pattern that cannot be generated by D0L systems. Theoretical Computer Science 83 (2), 1991, 205–218. (with M. J. M. de Boer and A. Lindenmayer)

75. Extensions of Gallai’s graph covering theorems for uniform hypergraphs. Journal of Combinatorial Theory, Ser. B 52 (1), 1991, 92–96.

76. Representations of relation algebras and patterns of colored triplets. “ Algebraic Logic ” (H. Andréka et al., eds.), North-Holland, 1991, 671–693.

77. Perfect graph decompositions. Graphs and Combinatorics 7 (1), 1991, 89–93.

78. Decompositions of regular bipartite graphs. Discrete Mathematics 89 (1), 1991, 17–27. (with M. S. Jacobson and M. Truszczyński)

79. Local constraints ensuring small representing sets. Journal of Combinatorial Theory, Ser. A 58 (1), 1991, 78–84. (with P. Erdős and A. Hajnal)

80. Dissections into equilateral triangles. Elemente der Mathematik 46 (6), 1991, 153–158.

81. Saturated r-uniform hypergraphs. Discrete Mathematics 98 (2), 1991, 95–104. (with P. Erdős and Z. Füredi)

82. Graph coloring problems with applications in algebraic logic. Acta Universitatis Carolinae, Mathematica et Physica 32 (2), 1991, 55–60.

83. Narrowness, pathwidth, and their application in natural language processing. Discrete Applied Mathematics 36 (1), 1992, 87–92. (with A. Kornai)

84. Rainbow subgraphs in properly edge-colored graphs. Random Structures and Algorithms 3 (2), 1992, 175–182. (with V. Rödl)

85. Graph coloring in linear time. Journal of Combinatorial Theory, Ser. B 55 (2), 1992, 236–243.

86. Precoloring extension. I. Interval graphs. Discrete Mathematics 100 (1–3), 1992, 267–279. (with M. Biró and M. Hujter)

Reprinted in: “ The Julius Petersen Graph Theory Centennial,” Topics in Discrete Mathematics 6, Elsevier, 1992.

87. Rado’s Selection Principle: applications to binary relations, graph and hypergraph colorings and partially ordered sets. Discrete Mathematics 103 (3), 1992, 301–312. (with M. Truszczyński)

88. Locally generated colourings of hexagonal cell division patterns; application to retinal cell differentiation. “ Lindenmayer Systems—Impacts on Theoretical Computer Science, Computer Graphics, and Developmental Biology ” (G. Rozenberg and A. Salomaa, eds.), Springer-Verlag, 1992, 333–350. (with A. Lindenmayer)

89. Irregular assignments and vertex distinguishing edge-colorings of graphs. Annals of Discrete Mathematics 52, 1992, 1–9. (with M. Aigner and E. Triesch)

90. Construction of non-isomorphic regular tournaments. Annals of Discrete Mathematics 52, 1992, 11–23. (with A. Astié-Vidal and V. Dugat)

91. Large partial parallel classes in Steiner systems. Annals of Discrete Mathematics 52, 1992, 545–548.

92. Decompositions into subhypergraphs with the Kőnig property. “ Sets, Graphs and Numbers ” (G. Halász et al., eds.) North-Holland, 1992, 705–711.

93. Perfect couples of graphs. Combinatorica 12 (2), 1992, 179–192. (with J. Körner and G. Simonyi)

94. Covering the cliques of a graph with vertices. Discrete Mathematics 108 (1–3), 1992, 279–289. (with P. Erdős and T. Gallai)

Reprinted in: “ Topological, Algebraical and Combinatorial Structures,” Topics in Discrete Mathematics 8, Elsevier, 1992.

95. Asymptotic growth of sparse saturated structures is locally determined. Discrete Mathematics 108 (1–3), 1992, 397–402.

Reprinted in: “ Topological, Algebraical and Combinatorial Structures,” Topics in Discrete Mathematics 8, Elsevier, 1992.

96. Small transversals in uniform hypergraphs. Siberian Advances in Mathematics 2 (1), 1992, 82–88. (with P. Erdős, D. Fon-Der-Flaass, and A. V. Kostochka)

97. Theorem proving through depth-first test. Acta Universitatis Carolinae, Mathematica et Physica 33 (2), 1992, 135–141.

98. Nearly uniform distribution of edges among k-subgraphs of a graph. Journal of Graph Theory 16 (6), 1992, 591–604. (with J. Širáň)

99. Domination properties and induced subgraphs. Discrete Mathematics 111 (1–3), 1993, 37–40. (with G. Bacsó)

100. Bounded degrees and prescribed distances in graphs. Discrete Mathematics 111 (1–3), 1993, 87–93. (with Y. Caro)

101. Helly property in finite set systems. Journal of Combinatorial Theory, Ser. A 62 (1), 1993, 1–14.

102. Complementary pairs of graphs orientable to line digraphs. Journal of Combinatorial Mathematics and Combinatorial Computing 13, 1993, 65–75. (with Z. Radosavljević and S. Simić)

103. Multipartite Turán problem for connected graphs and hypergraphs. Discrete Mathematics 112 (1–3), 1993, 199–206.

104. Precoloring extension. II. Graph classes related to bipartite graphs. Acta Mathematica Universitatis Comenianae 62 (1), 1993, 1–11. (with M. Hujter)

105. Algorithmic aspects of neighborhood numbers. SIAM Journal on Discrete Mathematics 6 (1), 1993, 24–29. (with G. J. Chang and M. Farber)

106. Rainbow subgraphs in edge-colorings of complete graphs. Annals of Discrete Mathematics 55, 1993, 81–88. (with P. Erdős)

107. One more occurrence of variables makes satisfiability jump from trivial to NP -complete. SIAM Journal on Computing 22 (1), 1993, 203–210. (with J. Kratochvíl and P. Savicky)

108. Graphs with no induced C_{4} and 2K_{2}. Discrete Mathematics 115 (1–3), 1993, 51–55. (with Z. Blázsik, M. Hujter, and A. Pluhár)

109. The number of maximal independent sets in triangle-free graphs. SIAM Journal on Discrete Mathematics 6 (2), 1993, 284–288. (with M. Hujter)

110. An upper bound on the number of cliques in a graph. Networks 23 (3), 1993, 207–210. (with M. Farber and M. Hujter)

111. Two graph-colouring games. Bulletin of the Australian Mathematical Society 48 (1), 1993, 141–149. (with F. Harary)

112. Menger-type theorems with restrictions on path lengths. Discrete Mathematics 120 (1–3), 1993, 161–174. (with L. Pyber)

113. Cross fertilisation of graph theory and aircraft maintenance scheduling. “ Thirty-Second Annual Symposium ” (G. Davison, ed.), AGIFORS (Airline Group of the International Federation of Operational Research Societies), 1993, 307–317. (with M. Biró and M. Hujter)

114. Linear-time approximation algorithms for the Max Cut problem. Combinatorics, Probability and Computing 2 (2), 1993, 201–210. (with N. V. Ngoc)

115. Large s-representable set systems with low maximum degree. Discrete Mathematics 122 (1–3), 1993, 205–217. (with P. Horák)

116. On k-local and k-mean colorings of graphs and hypergraphs. Quarterly Journal of Mathematics, Oxford (2) 44 (176), 1993, 385–398. (with Y. Caro)

117. Eulerian polynomial identities on matrix rings. Journal of Algebra 161 (1), 1993, 90–101. (with G. Révész and J. Szigeti)

118. On conjectures of Berge and Chvátal. Discrete Mathematics 124 (1–3), 1994, 79–86. (with M. Gionfriddo)

119. Bipartite subgraphs of triangle-free graphs. SIAM Journal on Discrete Mathematics 7 (2), 1994, 307–313. (with S. Poljak)

120. Algorithmic complexity of list colorings. Discrete Applied Mathematics 50 (3), 1994, 297– 302. (with J. Kratochvíl)

121. Monochromatic coverings and tree Ramsey numbers. Discrete Mathematics 125 (1–3), 1994, 377–384.

122. Inequalities for minimal covering sets in set systems of given rank. Discrete Applied Mathematics 51 (1–2), 1994, 187–195.

123. Largest size and union of Helly families. Discrete Mathematics 127 (1–3), 1994, 319–327.

124. Applications of the set-pair method in extremal hypergraph theory. “ Extremal Problems for Finite Sets ” (P. Frankl et al., eds.), Bolyai Society Mathematical Studies 3, János Bolyai Math. Soc., Budapest, 1994, 479–514.

125. Intersection dimensions of graph classes. Graphs and Combinatorics 10 (2), 1994, 159–168. (with J. Kratochvíl)

126. The spectrum of λ-times repeated blocks for T S(v, λ). Discrete Mathematics 129 (1–3), 1994, 159–166. (with S. Milici)

127. Solution to a problem of Speranza on M_{s}-colourings of directed graphs. Bolletino della Unione Matematica Italiana (7), 8-A, 1994, 283–286. (with M. Gionfriddo and S. Milici)

128. Packing problems in edge-colored graphs. Discrete Applied Mathematics 52 (3), 1994, 295–306. (with P. Hell and Y. Manoussakis)

129. Blocking sets in SQS(2v). Combinatorics, Probability and Computing 3, 1994, 77–86. (with M. Gionfriddo and S. Milici)

130. Characterization of (m, 1)-transitive and (3, 2)-transitive semi-complete directed graphs. Discrete Mathematics 135 (1–3), 1994, 335–347.

131. Perfect triangle families. Bulletin of the London Mathematical Society 26, 1994, 321–324.

132. Small n-dominating sets. Mathematica Pannonica 5 (2), 1994, 271–273.

133. Local and global average degree in graphs and multigraphs. Journal of Graph Theory 18 (7), 1994, 647–661. (with E. Bertram, P. Erdős, P. Horák, and J. Širáň)

134. The expected relative error of the polyhedral approximation of the max-cut problem. Operations Research Letters 16 (4), 1994, 191–198. (with S. Poljak)

135. Paths through fixed vertices in edge-colored graphs. Journal des Mathématiques, Informatique et Sciences Humaines 32 (127), 1994, 49–58. (with W.-S. Chou, Y. Manoussakis, O. Megalalaki, and M. Spyratos)

136. Rankings of graphs. “ Graph Theoretic Concepts in Computer Science ” (E. W. Mayr et al., eds.), Lecture Notes in Computer Science 903, Springer-Verlag, 1995, 292–304. (with H. Bodlaender, J. S. Deogun, K. Jansen, T. Kloks, D. Kratsch, and H. Müller)

137. Small representations of the relation algebra E_{n+1}(1, 2, 3). Algebra Universalis 33 (1), 1995, 136–139. (with P. Jipsen and R. Maddux)

138. 4-chromatic graphs with large odd girth. Discrete Mathematics 138 (1–3), 1995, 387–392. (with N. V. Ngoc)

139. The acyclic orientation game on random graphs. Random Structures and Algorithms 6 (2–3), 1995, 261–268. (with N. Alon)

140. How to make a random graph irregular. Random Structures and Algorithms 6 (2–3), 1995, 323–329.

141. Zero-sum block designs and graph labelings. Journal of Combinatorial Designs 3 (2), 1995, 89–99.

142. Vertex coverings of the edge set in a connected graph. “ Graph Theory, Combinatorics, and Applications ” (Y. Alavi and A. J. Schwenk, eds.), John Wiley & Sons 1995, 1179–1187. (with P. Erdős)

143. Searching for acyclic orientations of a graph. Discrete Mathematics 144 (1–3), 1995, 3–10. (with M. Aigner and E. Triesch)

144. Maximum cuts and large bipartite subgraphs. “ Combinatorial Optimization ” (W. Cook et al., eds.), DIMACS Series in Discrete Mathematics and Theoretical Computer Science 20, American Mathematical Society, 1995, 181–244. (with S. Poljak)

A somewhat shortened version of the preprint “ The Max-Cut problem — A survey.” Academia Sinica, Taipei, Taiwan, 1993, 91 pp.

145. Coverable graphs. Studia Scientiarum Mathematicarum Hungarica 30 (4), 1995, 329–344. (with S. Milici)

146. Degree sums and covering cycles. Journal of Graph Theory 20 (4), 1995, 419–422. (with H. Enomoto, A. Kaneko, and M. Kouider)

147. Stronger bounds for generalized degrees and Menger path systems. Discussiones Mathematicae – Graph Theory 15 (2), 1995, 167–177. (with R. J. Faudree)

148. Extremum and convergence of local average degrees in graphs. Congressus Numerantium 112, 1995, 191–198. (with S. Poljak and T. Szabó)

149. Cycle systems without 2-colorings. Journal of Combinatorial Designs 4 (2), 1996, 135–142. (with S. Milici)

150. Precoloring extension. III. Classes of perfect graphs. Combinatorics, Probability and Computing 5 (1), 1996, 35–56. (with M. Hujter)

151. On a conjecture of Erdős, Rubin and Taylor. Tatra Mountains Mathematical Publications 9, 1996, 69–82. (with M. Voigt)

152. Applications of the set-pair method in extremal problems, II., “ Combinatorics, Paul Erdős is Eighty ” (D. Miklós et al., eds.), Bolyai Society Mathematical Studies 2, János Bolyai Math. Soc., Budapest, 1996, 459–490.

153. Optimal routings in communication networks with linearly bounded capacity. Networks 28, 1996, 177–180. (with Y. Manoussakis)

154. Cycle-saturated graphs of minimum size. Discrete Mathematics 150 (1–3), 1996, 31–48. (with C. A. Barefoot, L. H. Clark, R. C. Entringer, T. D. Porter, and L. A. Székely)

155. Covering and independence in triangle structures. Discrete Mathematics 150 (1–3), 1996, 89–101. (with P. Erdős and T. Gallai)

156. The irregularity strength and cost of the union of cliques. Discrete Mathematics 150 (1–3), 179–186. (with S. Jendrol’ and M. Tkáč)

157. Cycles of given color patterns. Journal of Graph Theory 21 (2), 1996, 153–162. (with Y. Manoussakis and M. Spyratos)

158. Process structure driven control structure selection. IFAC World Congress 1996, Vol. M, 187–192. (with K. M. Hangos)

159. The forwarding index of directed networks. Discrete Applied Mathematics 68 (3), 1996, 279–291. (with Y. Manoussakis)

160. Ramsey-remainder. European Journal of Combinatorics 17 (6), 1996, 519–532. (with P. Erdős and P. Valtr)

161. Every 2-choosable graph is (2m, m)-choosable. Journal of Graph Theory 22 (3), 1996, 245–252. (with M. Voigt)

162. Computational aspects of graph theoretic methods in control. “ Computer-Intensive Methods in Control and Signal Processing—Can We Beat the Curse of Dimensionality? ” (L. Berec et al., eds.), 2nd European IEEE Workshop, Prague, Czech Republic, 1996, 25–36. (with K. M. Hangos)

163. Some hypergraph questions suggested by a Nordic Olympiad problem. Journal of Combinatorial Mathematics and Combinatorial Computing 22, 1996, 183–192. (with C. C. Rousseau)

164. Hamiltonian properties of Toeplitz graphs. Discrete Mathematics 159 (1–3), 1996, 69–81. (with R. van Dal, G. Tijssen, J. A. A. van der Veen, Ch. Zamfirescu, and T. Zamfirescu)

165. Lower bounds for a class of depth-two switching circuits. “ Matrices and Graphs—Theory and Applications to Economics ” (S. Camiz and S. Stefani, eds.), World Scientific Publ. Co., 1996, 7–18.

166. Minimal colorings for properly colored subgraphs. Graphs and Combinatorics 12 (4), 1996, 345–360. (with Y. Manoussakis, M. Spyratos, and M. Voigt)

167. Comparability graph augmentation for some multiprocessor scheduling problems. Discrete Applied Mathematics 72 (1–2), 1997, 71–84. (with P. Dell’Olmo and M. G. Speranza)

168. Solving systems of linear equations over Lie nilpotent rings. Linear and Multilinear Algebra 42, 1997, 43–51. (with J. Szigeti)

169. Efficiency and effectiveness of normal schedules on three dedicated processors. Discrete Mathematics 164, 1997, 67–79. (with P. Dell’Olmo and M. G. Speranza)

170. Choosability and fractional chromatic numbers. Discrete Mathematics 165/166, 1997, 31–38. (with N. Alon and M. Voigt)

171. Generalized colorings and avoidable orientations. Discussiones Mathematicae – Graph Theory 17 (1), 1997, 137–145. (with J. Szigeti)

172. The color cost of a caterpillar. Discrete Mathematics 174 (1–3), 1997, 125–130. (with M. Gionfriddo and F. Harary)

173. Upper chromatic number of Steiner triple and quadruple systems. Discrete Mathematics 174 (1–3), 1997, 247–259. (with L. Milazzo)

174. Switching distance graphs. Ars Combinatoria 47, 1997, 109–119. (with J. Gimbel and M. A. Henning)

175. A weighted generalization of Turán’s theorem. Journal of Graph Theory 25 (4), 1997, 267–275. (with J. A. Bondy)

176. List colorings and reducibility. Discrete Applied Mathematics 79 (1–3), 1997, 247–256. (with M. Voigt)

177. Graph colorings with local constraints — A survey. Discussiones Mathematicae – Graph Theory 17 (2), 1997, 161–228.

178. Dominating subgraphs of small diameter. Journal of Combinatorics, Information and System Sciences 22 (1), 1997, 51–62. (with G. Bacsó)

179. Semi on-line algorithms for the partition problem. Operations Research Letters 21 (5), 1997, 235–242. (with H. Kellerer, V. Kotov and M. G. Speranza)

180. Rankings of graphs. SIAM Journal on Discrete Mathematics 11 (1), 1998, 168–181. (with H. Bodlaender, J. S. Deogun, K. Jansen, T. Kloks, D. Kratsch, and H. Müller)

(Extended version of [136])

181. Strict colourings for classes of Steiner triple systems. Discrete Mathematics 182 (1–3), 1998, 233–243. (with L. Milazzo)

182. A 13/12 approximation algorithm for bin packing with extendable bins. Information Processing Letters 65 (5), 1998, 229–233. (with P. Dell’Olmo, H. Kellerer and M. G. Speranza)

183. On the approximation of finding a(nother) Hamiltonian cycle in cubic Hamiltonian graphs. Proceedings of STACS’98, Lecture Notes in Computer Science 1373, Springer-Verlag, 1998, 276–286. (with C. Bazgan and M. Santha)

184. Brooks-type theorems for choosability with separation. Journal of Graph Theory 27 (1), 1998, 43–49. (with J. Kratochvíl and M. Voigt)

185. Efficient approximation algorithms for the Subset-Sums Equality problem. Proceedings of ICALP’98 (K. G. Larsen, S. Skyum and G. Winskel, eds.), Lecture Notes in Computer Science 1443, Springer-Verlag, 1998, 387–396. (with C. Bazgan and M. Santha)

186. Complexity of choosing subsets from color sets. Discrete Mathematics 191 (1–3), 1998, 139–148. (with J. Kratochvíl and M. Voigt)

187. Rankings of directed graphs. Graph-Theoretic Concepts in Computer Science, Lecture Notes in Computer Science 1517, Springer-Verlag, 1998,114–123. (with J. Kratochvíl)

188. On magnitude networks. Congressus Numerantium 133, 1998, 75–84. (with J.-L. Fouquet, G. Hahn, and R. E. Woodrow)

189. Maximum cuts: Improvements and local algorithmic analogues of the Edwards–Erdős inequality. Discrete Mathematics 194 (1–3), 1999, 39–58. (with S. Bylka and A. Idzik)

190. A comparison of heuristics for scheduling multiprocessor dedicated tasks on three processors. Parallel Computing 25 (1), 1999, 49–61. (with A. K. Amoura, E. Bampis, and Y. Manoussakis)

191. On-line approximation algorithms for scheduling tasks on identical machines with extendable working time. Annals of Operations Research 86, 1999, 491–506. (with M. G. Speranza)

192. Chromatic spectrum is broken. 6th Twente Workshop on Graphs and Combinatorial Optimization (H. J. Broersma et al., eds.), University of Twente / Elsevier Preprint ISSN 0169-2690, 1999, 94–97. (with T. Jiang, D. Mubayi, V. I. Voloshin, and D. West)

193. Representation, constraints and complexity of graphs in the PLA spatial database model. Proceedings of the Jubilee International Conference (I. J. Rudas and A. L. Bencsik, eds.), Bánki Donát Polytechnic, Budapest, 1999, 57–60. (with Á. B. Novák)

194. On the approximation of finding a(nother) Hamiltonian cycle in cubic Hamiltonian graphs. Journal of Algorithms 31 (1), 1999, 249–268. (with C. Bazgan and M. Santha)

(Extended version of [183])

195. On light cycles in plane triangulations. Discrete Mathematics 197/198, 1999, 453–467. (with S. Jendrol’, T. Madaras, and R. Soták)

196. Fractional P-colourings and P-choice-ratio. Tatra Mountains Mathematical Publications 18, 1999, 69–77. (with P. Mihók and M. Voigt)

197. New trends in the theory of graph colorings: Choosability and list coloring. “ Contemporary Trends in Discrete Mathematics ” (R. L. Graham et al., eds.), DIMACS Series in Discrete Mathematics and Theoretical Computer Science 49, American Mathematical Society, 1999, 183–197. (with J. Kratochvíl and M. Voigt)

198. Process model structure simplification. Computers and Chemical Engineering Supplement 1999, S343–S346. (with K. M. Hangos)

199. Disjoint blocking sets in cycle systems. Discrete Mathematics 208/209, 1999, 451–462. (with S. Milici)

200. Objects reconstruction and its complexity problems in the PLA spatial database model. 3rd IEEE International Conference on Intelligent Engineering Systems, INES’99 (I. J. Rudas and L. Madarász, eds.), 1999, 219–223. (with Á. B. Novák)

201. Rankings of directed graphs. SIAM Journal on Discrete Mathematics 12 (3), 1999, 374– 384. (with J. Kratochvíl)

(Extended version of [187])

202. Remarks on the existence of uniquely partitionable planar graphs. Discussiones Mathematicae – Graph Theory 19 (2), 1999, 159–166. (with M. Borowiecki, P. Mihók, and M. Voigt)

203. On the complexity of bicoloring clique hypergraphs of graphs. (Extended Abstract) 11th Annual ACM–SIAM Conference on Discrete Algorithms, SODA 2000 , 40–41. (with J. Kratochvíl)

204. Uncolorable mixed hypergraphs. Discrete Applied Mathematics 99 (1–3), 2000, 209–227. (with V. I. Voloshin)

205. On-line rankings of graphs. Discrete Mathematics 212 (1–2), 2000, 141–147. (with I. Schiermeyer and M. Voigt)

206. Extremal bi-Helly families. Discrete Mathematics 213 (1–3), 2000, 321–331.

207. Sharpness of some intersection theorems. European Journal of Combinatorics 21 (7), 2000, 913–916. (with M. Hujter and L. Spissich)

208. Capelli polynomials, almost-permutation matrices and sparse Eulerian graphs. Discrete Mathematics 230 (1–3), 2001, 49–61. (with A. Lee, G. Révész, and J. Szigeti)

209. Heredity properties of connectedness in edge-coloured complete graphs. Discrete Mathematics 235 (1–3), 2001, 301–306. (with A. Idzik)

210. Oriented list colorings of graphs. Journal of Graph Theory 36 (4), 2001, 217–229. (with M. Voigt)

211. Ramsey numbers for tournaments. Theoretical Computer Science 263 (1–2), 2001, 75–85. (with Y. Manoussakis)

212. Optimal control structure selection for process systems. Computers and Chemical Engineering 25, 2001, 1521–1536. (with K. M. Hangos)

213. Unsolved combinatorial problems, Part I. BRICS Lecture Series, LS-01-1, viii+30 pp., May 2001. ISSN 1395-2048, http://www.brics.dk/LS/Ref/BRICS-LS-Ref/BRICS-LS-Ref.html

214. Complexity of coloring graphs without forbidden induced subgraphs. “ Graph-Theoretic Concepts in Computer Science ” (A. Brandstädt and V. B. Le, eds.), Lecture Notes in Computer Science 2204, Springer-Verlag, 2001, 254–262. (with D. Král’, J. Kratochvíl, and G. Woeginger)

215. Minimally non-preperfect graphs of small maximum degree. Graphs and Combinatorics 17 (4), 2001, 759–773. (with A. Wagler)

216. Structure simplification of dynamic process models. Journal of Process Control 12, 2002, 69–83. (with K. M. Hangos and A. Leitold)

217. Reconstruction of graphs and testing their properties in a relational spatial database. Computers and Mathematics with Applications 43 (10–11), 2002, 1391–1406. (with Á. B. Novák)

(Extended version of [193] and [200])

218. Uniquely colorable mixed hypergraphs. Discrete Mathematics 248 (1–3), 2002, 221–236. (with V. I. Voloshin and H. Zhou)

219. Efficient approximation algorithms for the Subset-Sums Equality problem. Journal of Computer and System Sciences 64 (2), 2002, 160–170. (with C. Bazgan and M. Santha)

Extended version of [185]

220. A colourful theory on graphs. ERCIM News Vol. 50, Special Issue ERCIMathematics, July 2002, 38–39.

221. Structural domination of graphs. Ars Combinatoria 63, 2002, 235–256. (with G. Bacsó)

222. A note on planar 5-list colouring : Non-extendability at distance 4. Discrete Mathematics 251 (1–3), 2002, 169–172. (with M. Voigt)

223. On the complexity of bicoloring clique hypergraphs of graphs. Journal of Algorithms 45 (1), 2002, 40–54. (with J. Kratochvíl)

(Extended version of [203])

224. The chromatic spectrum of mixed hypergraphs. Graphs and Combinatorics 18 (2), 2002, 309–318. (with T. Jiang, D. Mubayi, V. I. Voloshin, and D. West)

(Extended version of [192])

225. Domination in partitioned graphs. Discussiones Mathematicae – Graph Theory 22 (1), 2002, 199–210. (with P. D. Vestergaard)

226. Domination number of graphs with bounded diameter. Journal of Combinatorial Mathematics and Combinatorial Computing 40, 2002, 183–191. (with S. Mohammed Al-Yakoob)

227. The effect of modelling assumptions on the differential index of lumped process models. “ European Symposium on Computer Aided Process Engineering ” – ESCAPE 12 (J. Grievink and J. van Schijndel, eds.), Elsevier, 2002, 979–984. (with K. M. Hangos and G. Szederkényi)

228. Distance graphs with finite chromatic number. Journal of Combinatorial Theory, Ser. B 85 (1), 2002, 181–187. (with I. Z. Ruzsa and M. Voigt)

229. On the b-chromatic number of graphs. “ Graph-Theoretic Concepts in Computer Science ” (L. Kučera, ed.), Lecture Notes in Computer Science 2573, Springer-Verlag, 2002, 310–320. (with J. Kratochvíl and M. Voigt)

230. Representing directed and weighted graphs and their properties in a relational spatial database model. 6th IEEE International Confenrence on Intelligent Engineering Systems, INES 2002 (A. Lovrenčić and I. J. Rudas, eds.), 2002, 357–360. (with Á. B. Novák)

231. Strict colorings of Steiner triple and quadruple systems : a survey. Discrete Mathematics 261 (1–3), 2003, 399–411. (with L. Milazzo and V. I. Voloshin)

232. Unique colorings of bi-hypergraphs. Australasian Journal of Combinatorics 27, 2003, 33–45. (with G. Bacsó and V. I. Voloshin)

233. Linear-time algorithm for the edge-colorability of a graph with prescribed vertex types. Computer Science Journal of Moldova 11 (1) (31), 2003, 35–42. (with V. I. Voloshin)

234. Constructing Hamiltonian paths of some special graphs, using a relational spatial database model. 7th IEEE International Conference on Intelligent Engineering Systems, INES 2003. (with Á. B. Novák)

235. Computational aspects of connectionist interaction information retrieval. “ Mathematical/ Formal Methods in Information Retrieval,” 26th Annual International ACM SIGIR Conference, Toronto, Canada, 2003, 9 pages. (with S. Dominich)

236. Discovering spanning trees in a spatial relational model. IEEE International Conference on Computational Cybernetics, Siófok, Hungary, 2003, 43–46. (with Á. B. Novák and Zs. Ronkay)

237. Marking games and the oriented game chromatic number of partial k-trees. Graphs and Combinatorics 19 (1), 2003, 121–129. (with H. Kierstead)

238. Schütte’s tournament problem and intersecting families of sets. Combinatorics, Probability and Computing 12 (4), 2003, 359–364. (with M. Borowiecki, J. Grytczuk, and M. Hałuszczak)

239. Semi-on-line scheduling on two parallel processors with an upper bound on the items. Algorithmica 37 (4), 2003, 243–262. (with E. Angelelli and M. G. Speranza)

240. Local anti-Ramsey numbers of graphs. Combinatorics, Probability and Computing 12 (5–6), 2003, 495–511. (with M. Axenovich and T. Jiang)

241. On the existence and determination of satisfactory partitions in a graph. “ Algorithms and Computation ” (T. Ibaraki et al., eds.), Lecture Notes in Computer Science 2906, 2003, 444–453. (with C. Bazgan and D. Vanderpooten)

242. Spatiotemporal database models for attracting students to research. Proceedings of the International Conference in Memoriam John von Neumann (I. J. Rudas et al., eds.), Budapest, Hungary, 2003, 244–256. (with Á. B. Novák and P. Revesz)

243. Graph coloring. “ Handbook of Graph Theory ” (J. L. Gross and J. Yellen, eds.), Discrete Mathematics and Its Applications Vol. 25, CRC Press, 2004, 341–364.

244. Further topics in graph coloring. “ Handbook of Graph Theory ” (J. L. Gross and J. Yellen, eds.), Discrete Mathematics and Its Applications Vol. 25, CRC Press, 2004, 365–388.

245. The effect of model simplification assumptions on the differential index of lumped process models. Computers and Chemical Engineering 28 (1–2), 2004, 129–137. (with K. M. Hangos and G. Szederkényi)

(Extended version of [227])

Papers to appear

246. Scheduling groups of tasks with precedence constraints on three dedicated processors. Discrete Applied Mathematics (accepted) (with R. Mansini and M. G. Speranza)

247. Semi on-line scheduling on three processors with known sum of the tasks. Journal of Scheduling (accepted) (with E. Angelelli and M. G. Speranza)

248. Strong branchwidth and local transversals. Discrete Applied Mathematics (accepted)

249. Graphs without P_{5} and C_{5}. Discussiones Mathematicae – Graph Theory (accepted) (with G. Bacsó)

Printed on March 2, 2004